Assume that having a child with brown eyes or blue eyes are equally likely and that eye color is independent of the other children. What is the probability of having 5 children with brown eyes?
Answer by stanbon(75887)

Assume that having a child with brown eyes or blue eyes are equally likely and that eye color is independent of the other children. What is the probability of having 5 children with brown eyes?
---
P(blue) = P(brown) = 1/2
-----
P(5 brown) = (1/2)^5 = 1/32
--------------
Cheers,
Stan H.
--------------